National Distributors has $1,000 face value bonds outstanding with a market price of $1,013. The bonds pay interest semiannually, mature in 11 years, and have a yield to maturity of 6.87 percent. What is the current yield?

6.95 percent

7.61 percent

6.48 percent

7.39 percent

6.77 percent
Answer #1

use financial calculator to determine the copoun payments

PV=-1013

FV=1000

N=11*2=22 semi annual

1/Y=6.87/2=3.435

Click CPT

Click PMT=35.20 is semi annual payment

Annual coupon payment=35.20*2=70.40

current yield=70.40/1013=6.95%